# CALT-TH 2023-052

### 15 Dec 2023 v1

@@ 2.2 Graph Theory
Figure 3: An unbalanced network, where all the edges have capacity 1, that does not satisfy the nesting property. We have ${\texttt{S}}(ABC)=1$ and ${\texttt{S}}(C)=2$. The red highlighted edges denote the only cycle flow that maximizes the flux through $ABC$, and it does not simultaneously maximize the flux through $C$.
@@ 3.1 Entropy as Cycle Flows
Given a balanced network ${\mathcal{N}}=(V,E,c)$, we designate a subset of its vertices $\partial V=[{\sf{N}}+1]\subseteq V$, which we call boundary vertices, such that they each correspond to the party $A_{i}$ for $i\in[{\sf{N}}+1]$ in our system, with $A_{{\sf{N}}+1}$ being the purifier. We will then identify the entropy of the parties $\bigcup_{i\in{\underline{I}}}A_{i}$ to be the maximum magnitude of a multi-cycle flow through ${\underline{I}}\subseteq\partial V$, as defined in (2.18). In other words, recalling (2.18), we define
@@
Before we check various properties of this definition, it is useful to give some examples of how to associate an entropy vector to a balanced network. First, we claim that the prescription above describes the entropy vectors of all holographic states. To see why, notice that any undirected graph ${\mathcal{G}}=(V,E)$ with weighted edges can be represented as a balanced network such that the maximum flow out of any subset of vertices remain unchanged. For every edge $e\in E$ with capacity $c(e)$ connecting vertices $v_{1}$ and $v_{2}$, we can replace it with two directed edges such that one directed edge goes from $v_{1}$ to $v_{2}$ with capacity $c(e)$ and the other goes from $v_{2}$ to $v_{1}$ with the same capacity $c(e)$ (see Figure 4). This gives rise to a balanced network ${\mathcal{N}}$, and the flow out of any vertex stays the same as that of the original graph ${\mathcal{G}}$.
@@
All holographic entropy vectors can be represented using the max cycle flux of balanced networks given in (3.1).
@@
For this network, every vertex is a boundary vertex. Furthermore, given any subset of boundary vertices, the multi-cycle flow with maximum magnitude through any subset of boundary vertices is the minimal cycle flow ${\mathcal{C}}=(\{e_{1},e_{2},e_{3},e_{4}\};1)$, which is the clockwise flow with magnitude $1$. By our definition, the magnitude of this cycle flow through all subsets of boundary vertices is
@@
From the above example, we can easily construct a balanced network corresponding to a $n$-party GHZ state. Labeling the regions $A_{1},\ldots,A_{n}$, we depict this state in Figure 6. As in the 4-party GHZ network, every vertex is a boundary vertex, and the multi-cycle flow maximizing the magnitude through any subset of boundary vertices is again the clockwise cycle flow with magnitude $1$. Thus we find
@@
Notice that this also captures the notion of a general hypergraph edge explored in [17]. For instance, we can represent hypergraph 7 given in Figure 4 of [17] as a cycle flow in Figure 7 below. We simply replace each 4-vertex hyperedge with the 4-vertex cyclic network shown in Figure 5 and every undirected edge with the two directed edges shown in Figure 4. We leave it as an exercise for the reader to check that given our prescription (3.1), we indeed reproduce the entropy vector of the hypergraph. Furthermore, we prove below that in general, entropy vectors representable by hypergraphs can be captured by cycle flows.
$A$$B$$C$$D$$O$
@@
Figure 7: This balanced network reproduces the entropy vector represented by hypergraph 7 in Figure 4 of [17]. All edges have capacity $1$.
Proposition 4.
@@
All entropy vectors representable by hypergraphs can be represented using the prescription given in (3.1) on balanced networks.
Proof.
Given a hypergraph model, consider replacing every undirected edge by two oppositely directed edges shown in Figure 4, and every hyperedge involving $n$ vertices by a $n$-cyclic network like the one shown in Figure 6 with the same capacity for each edge. [^10] The choice of orientation for these directed edge cycles that replace hyperedges will not matter. Neither does the order of the boundary vertices around each cycle. For any subset of vertices ${\underline{I}}$, its entropy in the hypergraph model is given by the min cut of the hypergraph. We now construct a multi-cycle flow ${\mathfrak{C}}$ as follows. For every edge in the min cut, we will include in ${\mathfrak{C}}$ the minimal cycle that traverses the corresponding oppositely directed edges to a vertex in ${\underline{I}}^{c}$ with magnitude equal to the capacity of the min cut edge. If the min cut involves a hyperedge connecting $n$ vertices, then we include in ${\mathfrak{C}}$ the cycle flow through the corresponding $n$-cyclic network with magnitude saturating the capacity of the edges. The cycle flow we constructed has flux exactly equal to the min cut magnitude, and so the flux of the max cycle flow through ${\underline{I}}$ must be at least equal to the min cut magnitude of ${\underline{I}}$. However, the min cut separates the boundary vertices ${\underline{I}}$ from ${\underline{I}}^{c}$, which implies by our construction that the vertices in ${\underline{I}}$ are disconnected from those in ${\underline{I}}^{c}$ in the residual graph ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$. As there exists no more edges between ${\underline{I}}$ and ${\underline{I}}^{c}$ in the residual graph, we conclude ${\mathfrak{C}}$ must also be a max cycle flow, proving the flux of the max cycle flow through ${\underline{I}}$ corresponds to the min cut of ${\underline{I}}$ in the hypergraph. ∎
@@
As a more involved example, consider the balanced network given in Figure 8. This network is of particular interest as the entropic pattern it realizes cannot be captured by hypergraphs. In other words, it demonstrates that the proposed cycle flow prescription generalizes beyond the hypergraph models of [17]. The entropy vector that this network yields corresponds precisely to that of the stabilizer “graph state 15” of [20], which proved that the entanglement structure of such a stabilizer state is unattainable by hypergraphs. It was later shown in [18] that the entropy vector of graph state 15 can be obtained by providing a min cut model on topological links while still realizing all hypergraph entropies, albeit obscuring any potentially dual flow perspective. The topological character of the models of [18], and their realization of graph state $15$, served as motivation for our construction of the state using directed networks in Figure 8, as well for our cycle flow prescription for computing entropy in general. Although proving an equivalence between our prescription and that of [18] at the level of entropy cones is beyond the scope of this paper, the fact that we are able to reproduce the subsystem entropies of graph state $15$ does demonstrate that cycle flows share key topological features with link models.
@@
Figure 8: This balanced network reproduces the entropy vector represented by “graph state 15” in Figure 1 of [20]. Its associated entropy vector is $(11111;2222212212;2222222222;22221;1)$, where we have ordered the components in lexicographic order.
@@ 3.2 Subadditivity
The entropy defined via cycle flows given in (3.1) satisfies subadditivity.
@@
Consider a graph with boundary vertices $\partial V={\underline{I}}\cup{\underline{J}}\cup{\underline{K}}$, where ${\underline{I}},{\underline{J}}$, and ${\underline{K}}$ are mutually disjoint. We will prove
@@
To compute ${\texttt{S}}({\underline{I}}{\underline{J}})$, let ${\mathfrak{C}}({\underline{J}},{\underline{I}})$ be a minimal multi-cycle flow that maximizes the magnitude through both ${\underline{J}}$ and ${\underline{I}}{\underline{J}}$ simultaneously; such a multi-cycle flow exists by Theorem 1. It then follows
@@ 3.3 Strong Subadditivity
The entropy defined via cycle flows given in (3.1) satisfies strong subadditivity.
@@
Using (3.11), we have
@@ 4 Discussion
In the generalization that we proposed in this paper, we do not have such a clean geometric picture. Rather than generalizing the notion of graphs to undirected hypergraphs and links, as was previously explored in [17, 18], we show that it suffices to consider max flows on directed graphs. However, rather than considering traditional flows with sources and sinks, we need to consider cycle flows, as this was crucial to ensure that purification symmetry is manifest. Utilizing cycle flows, we are able to show that our prescription for computing entanglement entropy satisfies both SA and SSA. Furthermore, we straightforwardly showed that both the conventional graph model representing holographic entropy states and the hypergraph model are special cases of our cycle flow model. Importantly, the entropy vector associated to the stabilizer state that could not be reproduced by the hypergraph model has a very simple realization in our cycle flow model (see Figure 8).
@@
We would still like to obtain some form of geometric interpretation for the cycle flow model. The usual intuition that undirected edges capture geometric connectivity seems subtle to extend to directed edges. However, the work of [34] did successfully provide a geometric picture for the GHZ-like entanglement that the hypergraphs of [17] capture, even though hyperedges are at first glance not geometric in the sense standard edges are. [^11] Indeed, what geometrizes hyperedges are patched Euclidean geometries which meet at a space of higher codimension. As such, these are non-locally Euclidean spaces (thus not manifolds). We thank Tom Hartman for pointing out this geometric interpretation of GHZ-like entropies. Perhaps one particularly tantalizing clue is that the entropies of GHZ states involving $n$ parties are represented simply by a directed loop involving $n$ vertices in our cycle flow model.
@@
Furthermore, in the conventional “bit thread” model involving max flows, one only cared about the thread configuration making up the final multicommodity flow, and not the individual flows. However, as is evidenced by Figure 1, for us the simple cycles making up a multi-cycle flow are actually important, and two multi-cycle flows with exactly the same flow along each edge actually correspond to different entropy vectors depending on their respective simple cycles composing them. Additionally, these individual cycles are directed. For an undirected graph, we can transform it into a directed graph by taking every undirected edge and replacing it with two oppositely directed edges, as is illustrated in Figure 4, and the cycle flow is obtained from the max flow traversing some set of edges and coming back along the corresponding oppositely directed edges. However, for general balanced graphs, the path we take “going there” can be different from the one we take “coming back.” It would be both interesting and worthwhile to understand the physical significance of these minimal cycle flows.
@@
Lastly, it is important to point out that we have only proved necessary, but not sufficient, conditions for the cycle flow model to describe quantum entropy vectors. It is possible that there exist entropy vectors that obey all instances of SA and SSA, and yet are not realizable (or approximated arbitrarily well) by any quantum state. In other words, if we define an entropy cone given by the entropy vectors realizable by cycle flows, it is not clear whether or not this cone coincides with the quantum entropy cone, contains the quantum entropy cone, or neither. We leave the complete exploration of the interplay between the entropy vectors realizable using our cycle flow model and the quantum entropy cone to future work.
@@ Acknowledgements
We would like to thank Nina Anikeeva, Newton Cheng, Tom Hartman, Matthew Headrick, Oliver Kosut, Shinsei Ryu, Vincent Su, Brian Swingle, and Tadashi Takayanagi for useful conversations. We would especially like to thank Ning Bao, Sebastian Fischetti, and William Munizzi, who were involved in earlier iterations of this work. T.H. is supported by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021-2817, the U.S. Department of Energy grant DE-SC0011632, and the Walter Burke Institute for Theoretical Physics. S.H-C. is supported by the U.S. Department of Energy Award DE-SC0021886. C.K. is supported by the U.S. Department of Energy under grant number DE-SC0019470 and by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021-2818.
@@ Appendix A Remarks on the Dual Program
The max cycle flow prescription we proposed in (3.1) for computing entanglement entropy is a linear program. Consequently, there exists a dual program. As we shall see, the dual program cannot be simply described via a min edge cut prescription; nevertheless, we will find it provides a complementary approach to the max flow prescription. In Section A.1, we will set up the linear program and its dual. Next, in Section A.2, we explain the analogous notion of a vertex cut. Lastly, we demonstrate in Section A.3 how to use the dual program to provide alternative proofs of SA. We assume reasonable familiarity with convex duality, and refer the reader to [35] for an excellent exposition on this subject.

### 12 Oct 2024 v3

@@ 2.2 Graph Theory
Figure 4: A balanced network, where all edges have capacity 1, that does not satisfy the nesting property. We have ${\texttt{S}}(BC)={\texttt{S}}(ABCD)=3$. Each colored cycle denotes a simple cycle in the only multi-cycle flow that maximizes the flux through $ABCD$, and it does not simultaneously maximize the flux through $BC$. Note that this directed network is planar (although we did not draw it in a manifestly planar manner for aesthetic reasons), so planarity of the network does not imply nesting.
@@ 3.1 Entropy as Cycle Flows
Given a nesting balanced network ${\mathcal{N}}=(V,E,c)$, we designate a subset of its vertices $\partial V=[{\sf{N}}+1]\subseteq V$, which we call boundary vertices, such that they each correspond to the party $A_{i}$ for $i\in[{\sf{N}}+1]$ in our system, with $A_{{\sf{N}}+1}$ being the purifier. We will then identify the entropy of the parties $\bigcup_{i\in{\underline{I}}}A_{i}$ to be the maximum magnitude of a multi-cycle flow through ${\underline{I}}\subseteq\partial V$, as defined in (2.18). In other words, recalling (2.18), we define
@@
Before we check various properties of this definition, it is useful to give some examples of how to associate an entropy vector to a nesting balanced network. First, we claim that the prescription above describes the entropy vectors of all holographic states. To see why, notice that any undirected graph ${\mathcal{G}}=(V,E)$ with weighted edges can be represented as a balanced network such that the maximum flow out of any subset of vertices remain unchanged. For every edge $e\in E$ with capacity $c(e)$ connecting vertices $v_{1}$ and $v_{2}$, we can replace it with two directed edges such that one directed edge goes from $v_{1}$ to $v_{2}$ with capacity $c(e)$ and the other goes from $v_{2}$ to $v_{1}$ with the same capacity $c(e)$ (see Figure 5). The nesting property is then inherited from the nesting property of conventional flows. This gives rise to a nesting balanced network ${\mathcal{N}}$, and the flow out of any vertex stays the same as that of the original graph ${\mathcal{G}}$.
@@
All holographic entropy vectors can be represented using the max cycle flux of nesting balanced networks given in (3.1).
@@
For this network, every vertex is a boundary vertex. Furthermore, given any subset of boundary vertices, the multi-cycle flow with maximum magnitude through any subset of boundary vertices is the minimal cycle flow ${\mathcal{C}}=(\{e_{1},e_{2},e_{3},e_{4}\};1)$, which is the clockwise flow with magnitude $1$. Thus, this network trivially satisfies nesting. By our definition, the magnitude of this cycle flow through all subsets of boundary vertices is
@@
From the above example, we can easily construct a nesting balanced network corresponding to a $n$-party GHZ state. Labeling the regions $A_{1},\ldots,A_{n}$, we depict this state in Figure 7. As in the 4-party GHZ network, every vertex is a boundary vertex, and the multi-cycle flow maximizing the magnitude through any subset of boundary vertices is again the clockwise cycle flow with magnitude $1$. Thus we find
@@
Notice that this also captures the notion of a general hypergraph edge explored in [17]. For instance, we can represent hypergraph 7 given in Figure 4 of [17] as a cycle flow in Figure 8 below. We simply replace each 4-vertex hyperedge with the 4-vertex cyclic network shown in Figure 6 and every undirected edge with the two directed edges shown in Figure 5. We leave it as an exercise for the reader to check that given our prescription (3.1), we indeed reproduce the entropy vector of the hypergraph. Furthermore, we conjecture below that in general, entropy vectors representable by hypergraphs can be captured by cycle flows.
$A$$B$$C$$D$$O$22Refer to caption
@@
Figure 8: The network on the right reproduces the entropy vector represented by hypergraph 7 in Figure 4 of [17], which we reproduced on the left for convenience. All edges have capacity $1$ unless otherwise labeled.
Conjecture 1.
@@
All entropy vectors representable by hypergraphs can be represented using the prescription given in (3.1) on nesting balanced networks.
To understand the rationale behind this conjecture, note that given any hypergraph model, we can replace every undirected edge by two oppositely directed edges shown in Figure 5, and every hyperedge involving $n$ vertices by a $n$-cyclic network like the one shown in Figure 7 with the same capacity for each edge. There is still an ambiguity in the ordering of the vertices and the orientation of each such directed edge cycles replacing the hyperedges, but in many examples considered, we were able to find an orientation where the entropy vector represented by the hypergraph is reproduced by our directed network. Nevertheless, such a naive argument cannot be true in general, as the balanced network considered in Figure 4 does not have a hypergraph realization (since hypergraphs obey SSA), even though our above prescription for replacing directed edges with edges and hyperedges would suggest otherwise. [^13] We thank Massimiliano Rota for pointing this out. We therefore leave proving the conjecture an open question.
@@
Although we have not proved the above conjecture, the ability for directed networks to represent entropy vectors of quantum states not representable by hypergraphs is illustrated by the balanced network given in Figure 9. This network is of particular interest as the entropic pattern it realizes cannot be captured by hypergraphs. In other words, it demonstrates that the proposed cycle flow prescription generalizes beyond the hypergraph models of [17] (assuming 1). The entropy vector that this network yields corresponds precisely to that of the stabilizer “graph state 15” of [20], which proved that the entanglement structure of such a stabilizer state is unattainable by hypergraphs. It was later shown in [18] that the entropy vector of graph state 15 can be obtained by providing a min cut model on topological links while still realizing all hypergraph entropies, albeit obscuring any potentially dual flow perspective. The topological character of the models of [18], and their realization of graph state $15$, served as motivation for our construction of the state using directed networks in Figure 9, as well for our cycle flow prescription for computing entropy in general. Although proving an equivalence between our prescription and that of [18] at the level of entropy cones is beyond the scope of this paper, the fact that we are able to reproduce the subsystem entropies of graph state $15$ does demonstrate that cycle flows share key topological features with link models.
@@
Figure 9: This network reproduces the entropy vector represented by “graph state 15” in Figure 1 of [20]. Its associated entropy vector is $(11111;1222222221;2222222222;22122;1)$, where we have ordered the components in lexicographic order.
@@ 3.2 Subadditivity
The entropy defined via cycle flows on a nesting balanced network given in (3.1) satisfies subadditivity.
@@
Consider a network with boundary vertices $\partial V={\underline{I}}\cup{\underline{J}}\cup{\underline{K}}$, where ${\underline{I}},{\underline{J}}$, and ${\underline{K}}$ are mutually disjoint. We will prove
@@
To compute ${\texttt{S}}({\underline{I}}{\underline{J}})$, let ${\mathfrak{C}}({\underline{J}},{\underline{I}})$ be a minimal multi-cycle flow that maximizes the magnitude through both ${\underline{J}}$ and ${\underline{I}}{\underline{J}}$ simultaneously; such a multi-cycle flow exists by the definition of a nesting balanced network (see 9). It then follows
@@ 3.3 Strong Subadditivity
The entropy defined via cycle flows on a nesting balanced network given in (3.1) satisfies strong subadditivity.
@@
As our network allows for nested flows, we can use (3.11) to obtain
@@ 4 Discussion
In the generalization that we proposed in this paper, we do not have such a clean geometric picture. Instead of generalizing the notion of graphs to undirected hypergraphs and links, as was previously explored in [17, 18], we show that it suffices to consider max flows on directed graphs. However, rather than considering traditional flows with sources and sinks, we need to consider cycle flows, as this was crucial to ensure that purification symmetry is manifest. Utilizing cycle flows, we are able to show that our prescription for computing entanglement entropy satisfies both SA and SSA. Furthermore, we straightforwardly showed that the conventional graph model representing holographic entropy states is a special case of our cycle flow model, and similarly conjectured that the hypergraph model is a special case of our cycle flow model as well. [^14] It would be interesting to also attempt to connect our cycle flow model to the hyperthreads considered in [34, 35]. Importantly, the entropy vector associated to the stabilizer state that could not be reproduced by the hypergraph model has a very simple realization in our cycle flow model (see Figure 9).
@@
We would still like to obtain some form of geometric interpretation for the cycle flow model. The usual intuition that undirected edges capture geometric connectivity seems subtle to extend to directed edges. However, the work of [36] did successfully provide a geometric picture for the GHZ-like entanglement that the hypergraphs of [17] capture, even though hyperedges are at first glance not geometric in the sense standard edges are. [^15] Indeed, what geometrizes hyperedges are patched Euclidean geometries which meet at a space of higher codimension. As such, these are non-locally Euclidean spaces (thus not manifolds). We thank Tom Hartman for pointing out this geometric interpretation of GHZ-like entropies. Perhaps one particularly tantalizing clue is that the entropies of GHZ states involving $n$ parties are represented simply by a directed loop involving $n$ vertices in our cycle flow model. This clue, however, also presents a new mystery: the ordering of the vertices in the directed loop is arbitrary. Any ordering produces the same entropy vector. We regard this arbitrariness as a feature – since many states can have the same GHZ-like entropy vector, perhaps our directed networks actually contain information beyond the entropy vector.
@@
Furthermore, in the conventional “bit thread” model involving max flows, one only cared about the thread configuration making up the final multi-commodity flow, and not the individual flows. However, as is evidenced by Figure 1, for us the simple cycles making up a multi-cycle flow are actually important, and two multi-cycle flows with exactly the same flow along each edge actually correspond to different entropy vectors depending on their respective simple cycles composing them. Additionally, the individual simple cycles composing the multi-cycle flows are directed. For an undirected graph, we can transform it into a directed graph by taking every undirected edge and replacing it with two oppositely directed edges, as is illustrated in Figure 5, and the cycle flow is obtained from the max flow traversing some set of edges and coming back along the corresponding oppositely directed edges. However, for general balanced networks, the path we take “going there” can be different from the one we take “coming back.” It would be both interesting and worthwhile to understand the physical significance of these minimal cycle flows.
Indeed, one possible avenue towards better understanding the cycle flows is an improved understanding of the dual program. Because the primal variables are the flows along the cycles (and not the magnitude of the flows along edges), the dual variables, which are variables for the edges, do not result in a simple edge-cut prescription. Although we were successful in setting up the dual program in Appendix A and used it to provide alternative proofs of SA, we leave using it to find an alternative proof for SSA, as well as more generally a construction of such an edge-cut protocol, for future work. Indeed, the dual program set up in Appendix A does not require any restrictions coming from considering only nesting balanced networks, so it has the added benefit of proving all balanced networks satisfy SA. We hope that this will also inform us whether there is a cleaner way to categorize the types of balanced networks that satisfy nesting besides Theorem 1, since while Theorem 1 gives necessary and sufficient conditions for nesting to be satisfied, it is not straightforward to check that a random balanced network satisfies nesting.
@@
Lastly, it is important to point out that we have only proved necessary, but not sufficient, conditions for the cycle flow model to describe quantum entropy vectors. It is possible that there exist entropy vectors that obey all instances of SA and SSA, and yet are not realizable (or approximated arbitrarily well) by any quantum state. In other words, if we define an entropy cone given by the entropy vectors realizable by cycle flows, it is not clear whether or not this cone coincides with the quantum entropy cone, contains the quantum entropy cone, or neither. The relationship between the entropy cone defined by cycle flows and that defined by either hypergraphs [17] or links [18] is also unknown; although we conjectured in 1 that the hypergraph cone is a subset of the cycle flow cones, proving this is an important next step. We leave the complete exploration of the interplay between the entropy vectors realizable using our cycle flow model and the quantum entropy cone to future work.
@@ Acknowledgements
We would like to thank Nina Anikeeva, Newton Cheng, Tom Hartman, Matthew Headrick, Oliver Kosut, Shinsei Ryu, Vincent Su, Brian Swingle, and Tadashi Takayanagi for helpful conversations, and especially Massimiliano Rota for useful feedback on the preprint. We would further like to thank Ning Bao, Sebastian Fischetti, and William Munizzi, who were involved in earlier iterations of this work. T.H. would like to dedicate this to the memory of his father Yuanzhan He, whose unwavering support made it possible for T.H. to be a better physicist and, more importantly, a better person. T.H. is supported by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021-2817, the U.S. Department of Energy grant DE-SC0011632, and the Walter Burke Institute for Theoretical Physics. S.H.-C. is supported by the U.S. Department of Energy Award DE-SC0021886. C.K. is supported by the U.S. Department of Energy under grant number DE-SC0019470 and by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021-2818.
@@ Appendix A Remarks on the Dual Program
The max cycle flow prescription we proposed in (3.1) for computing entanglement entropy is a linear program. Consequently, there exists a dual program. As we shall see, the dual program cannot be simply described via a min edge cut prescription; nevertheless, we will find it provides a complementary approach to the max flow prescription. In Section A.1, we will set up the linear program and its dual. Next, in Section A.2, we explain the analogous notion of a vertex cut. Lastly, we demonstrate in Section A.3 how to use the dual program to provide alternative proofs of SA. Because our dual program does not need the added assumption of a nesting balanced network, our proof works for balanced networks in general. We assume reasonable familiarity with convex duality, and refer the reader to [37] for an excellent exposition on this subject.